# References

### v1

@@ I Introduction
Many applications require autonomous robots to perform tasks in complex, cluttered, dynamic, or unknown environments. In order for such a robot to adapt to new tasks and environments, an expert is often required to understand the required task logic and translate it into an implementation based on the robot’s capabilities. This severely limits speed and flexibility for robot deployment. LLMs promise to relax this requirement while allowing for intelligent replanning during task execution. To this end, we focus on the challenging field of service robotics, requiring robots to interact with non-expert operators in open-ended, highly unstructured environments designed for and shared with humans.
@@ II Related Works
The use of LLMs for high-level robot task planning and for low-level execution control are hot research topics [abs-2311-07226, abs-2401-04334].
@@ III-A Environment
In order to create an environment that allows an accurate comparison of the effect of prompt engineering techniques, we adopt the setting found at RoboCup@Home [rulebook_2024], which is a major international competition that focuses on domestic service robots.
The arena is designed to replicate a typical domestic dwelling with several rooms: A study room, a parlor, a kitchen, and a bedroom. All tasks start with both robot and operator located in the parlor. Each room contains multiple objects, including common food and household items, with which the robot can interact with.
@@ III-C Tasks
success rate mean time success time success time success time
Baseline 0.02 6.55s 0.00 9.07s 0.00 7.83s 0.00 10.47s
AF 0.22 3.99s 0.00 6.41s 0.00 5.23s 0.04 11.99
AF + EiP 1.00 5.57s 0.38 5.43s 0.20 9.45s 0.02 13.79s
AF + CoT 0.46 5.66s 0.02 5.61s 0.00 6.65s 0.00 8.25s
AF + CoT + EiP 1.00 6.76s 0.44 7.45s 0.16 11.47s 0.02 17.46s
AF + ReAct + EiP 1.00 14.88s 0.44 13.76s 0.10 22.67s 0.10 25.86s
AF + StD 0.82 4.93s 0.00 10.12s 0.00 8.96s 0.00 9.58s
AF + CoT + EiP + StD 1.00 8.93s 0.36 14.03s 0.00 16.40s 0.00 20.85s
AF + ReAct + EiP + StD 1.00 20.69s 0.72 41.71s 0.14 28.37s 0.00 68.19s
@@ III-C1 Fetch
The Fetch task requires the robot to pick up an object from another location and return it to the user. Since it can be solved with a simple sequence of function calls, it mainly tests whether the LLM, in combination with the applied prompt engineering technique, is able to comprehend the scenario, use the function-calling feature, and generate valid JSON appropriate to solve the task. An example of this task is "Please get me a pen from the study".
@@ III-C2 Conditional
The Conditional task extends the Fetch task by adding the requirement to gather knowledge about the environment during the execution of the task and to adapt to it. In particular, the robot is asked to search for an object and return one of two other objects to the user. The choice of the latter depends on the first object found. This task therefore tests whether the LLM is capable of branching logic. An example of this task is "Check if there is a spoon in the kitchen. If you find one, bring me a pen from the study. If not, bring me a comb from the bedroom".
@@ III-C3 Equals
The Equals task requires the robot to make a numerical observation and repeatedly retrieve an object according to that observation. It therefore tests basic mathematical ability and requires the execution of a logical loop. This results in a long task that requires many function calls, challenging the LLM to maintain focus on accomplishing the task objective. An example of this task is "For every orange in the kitchen, move a fork from the kitchen to the parlor".
@@ III-C4 Distribute
In the Distribute task, a given object must be distributed so that every location in the simulation contains at least one instance. The LLM must therefore not only visit all locations, but also make numerous decisions regarding the movement of objects and keep track of their locations. An example of this task is "Please distribute the pens evenly so that each location contains at least one pen. You can start with the pens in the study".
@@ IV-A LLM Variants
We focus on evaluating GPT [Fewshot] model variants, due to the function-calling feature of the corresponding API, and their top-of-the-line results on benchmarks regarding this feature [Leaderboard]. In particular, we use GPT-3.5-Turbo-0125 and GPT-4-0125-Preview, referred to as GPT 4 Turbo. We evaluate each condition with respect to tasks and prompt engineering techniques with 50 repetitions for GPT 3.5 Turbo and 20 repetitions for GPT 4 Turbo due to budgetary considerations. For both model variants and all tasks we use a temperature setting of $\tau=0$ to elicit factual responses. All other model parameters are set to the default values suggested by OpenAI.
@@ IV-B Prompt Engineering Techniques
Our simulation procedure allows the robot to execute tasks immediately after providing the LLM with the task description in an initial prompt along with the function definitions. Task completion then proceeds as a dialog of function calls generated by the LLM and corresponding function responses generated by the simulation. This approach serves as a baseline for investigating and comparing the effects of various prompt engineering techniques on task completion and time required. We focus on a selection of popular and applicable prompt engineering techniques for task planning and execution in the service robotics domain. Figure 2 gives an overview of this selection.
@@ IV-B2 Chain-of-Thought (CoT)
We integrate CoT [CoT] prompting into our framework by asking the LLM to first provide a textual step-by-step plan detailing how it intends to complete the task given the available functions from the function library. Only after such a plan has been generated do we ask the LLM to call the appropriate functions based on that plan. If the task execution turns out to require many steps, we repeat this planning step every 15 function calls, to ensure that the LLM is always operating according to a plan in case the initial plan is not detailed enough or incomplete.
@@ IV-B4 Example in Prompt
success rate mean time success time success time success time
Baseline 1.00 11.57s 0.90 14.72s 0.85 23.65s 0.45 37.38s
AF 1.00 10.69s 0.65 12.06s 0.60 22.68s 0.90 33.64s
AF + EiP 1.00 11.21s 1.00 16.03s 0.70 21.57s 0.65 31.97s
AF + CoT 1.00 15.14s 0.70 18.95s 0.70 33.03s 0.20 42.32s
AF + CoT + EiP 1.00 15.07s 1.00 20.69s 0.65 26.26s 0.80 36.89s
AF + ReAct + EiP 1.00 31.05s 1.00 31.44s 0.95 66.45s 0.70 64.25s
AF + StD 1.00 10.45s 0.90 12.63s 0.65 32.05s 0.90 29.57s
AF + CoT + EiP + StD 1.00 15.52s 1.00 20.63s 0.80 28.80s 0.75 36.44s
AF + ReAct + EiP + StD 1.00 29.38s 1.00 32.86s 1.00 72.34s 0.50 72.25s
@@ V Results
We find that more complex prompt engineering techniques do not necessarily lead to higher success rates. On the Equals and Distribute tasks with GPT 3.5 Turbo, the techniques without state descriptions outperform the corresponding technique with state descriptions. However, state descriptions do not always have a negative impact. For the GPT 3.5 Turbo Conditional task, the addition of state descriptions (AF + ReAct + EiP + StD) achieves the highest success rate of 72%. We also observe that using only adaptive functions achieves the best success rate with GPT 4 Turbo for the Distribute task, while achieving the worst success rate for the Conditional and Equals tasks. There’s also no clear favorite between CoT and ReAct, as the success rates vary depending on the task, model variant, and other prompt engineering techniques used in conjunction.
Due to the additional prompts used in all prompting techniques besides the adaptive functions, the measured task completion time depends on the the technique used. In general, we observe that CoT and even more so ReAct show a longer runtime due to the relatively large number of tokens generated by the LLM for planning and reasoning compared to function calls. Using state descriptions increases the time even more for the same reason. Even though the adaptive functions do not utilize additional prompts, we can see a slight reduction in time compared to the baseline, because the LLM is less likely to call functions that result in failures. In a real application, the response time of an LLM can significantly increase the execution time of a task, although functions that require considerable time themselves provide leeway to mask this problem.

### v2

@@ I Introduction
Many applications require autonomous robots to perform tasks in complex, cluttered, dynamic, or unknown environments. In order for such a robot to adapt to new tasks and environments, an expert is often required to understand the task logic and translate it into an implementation based on the robot’s capabilities. This severely limits speed and flexibility for robot deployment. LLMs promise to relax this requirement while allowing for intelligent replanning during task execution. To this end, we focus on the challenging field of service robotics [StuecklerSB:Cosero16, Memmesheimer:Winner2024], requiring robots to interact with non-expert operators in open-ended, highly unstructured environments designed for and shared with humans.
@@ II Related Works
Using LLMs for high-level robot task planning and for low-level execution control are hot research topics [abs-2311-07226, abs-2401-04334].
@@ III-A Environment
In order to create an environment that allows an accurate comparison of the effect of prompt engineering techniques, we adopt the setting found at RoboCup@Home [Stueckler:RAM2012, rulebook_2024], which is a major international competition that focuses on domestic service robots.
The arena is designed to replicate a typical domestic dwelling with several rooms: A study room, a parlor, a kitchen, and a bedroom. All tasks start with both robot and operator located in the parlor. Each room contains multiple objects, including common food and household items, with which the robot can interact.
@@ III-C Tasks
success rate mean time [s] success time [s] success time [s] success time [s]
Baseline 0.02 6.55 0.00 9.07 0.00 7.83 0.00 10.47
AF 0.22 3.99 0.00 6.41 0.00 5.23 0.04 11.99
AF + EiP 1.00 5.57 0.38 5.43 0.20 9.45 0.02 13.79
AF + CoT 0.46 5.66 0.02 5.61 0.00 6.65 0.00 8.25
AF + CoT + EiP 1.00 6.76 0.44 7.45 0.16 11.47 0.02 17.46
AF + ReAct + EiP 1.00 14.88 0.44 13.76 0.10 22.67 0.10 25.86
AF + StD 0.82 4.93 0.00 10.12 0.00 8.96 0.00 9.58
AF + CoT + EiP + StD 1.00 8.93 0.36 14.03 0.00 16.40 0.00 20.85
AF + ReAct + EiP + StD 1.00 20.69 0.72 41.71 0.14 28.37 0.00 68.19
@@ III-C1 Fetch
The Fetch task requires the robot to pick up an object from another location and to return it to the user. Since Fetch can be solved with a simple sequence of function calls, it mainly tests whether the LLM, in combination with the applied prompt engineering technique, is able to comprehend the scenario, use the function-calling feature, and generate valid JSON appropriate to solve the task. An example of Fetch is: "Please get me a pen from the study".
@@ III-C2 Conditional
The Conditional task extends Fetch by adding the requirement to gather knowledge about the environment during task execution and to adapt its behavior accordingly. In particular, the robot is asked to search for an object and return one of two other objects to the user. The choice of the latter depends on the first object found. This task therefore tests whether the LLM is capable of branching logic. An example of Conditional is: "Check if there is a spoon in the kitchen. If you find one, bring me a pen from the study. If not, bring me a comb from the bedroom".
@@ III-C3 Equals
The Equals task requires the robot to make a numerical observation and repeatedly retrieve an object according to that observation. It therefore tests basic mathematical ability and requires the execution of a logical loop. This results in a long task that requires many function calls, challenging the LLM to maintain focus on accomplishing the task objective. An example of Equals is: "For every orange in the kitchen, move a fork from the kitchen to the parlor".
@@ III-C4 Distribute
In the Distribute task, a given object must be distributed so that every location in the simulation contains at least one instance. The LLM must therefore not only visit all locations, but also make numerous decisions regarding the movement of objects and keep track of their locations. An example of Distribute is: "Please distribute the pens evenly so that each location contains at least one pen. You can start with the pens in the study".
@@ IV-A LLM Variants
We focus on evaluating GPT [Fewshot] model variants, due to the function-calling feature of the corresponding API, and their top-of-the-line results on benchmarks regarding this feature [Leaderboard]. In particular, we use GPT-3.5-Turbo-0125 and GPT-4-0125-Preview, referred to as GPT 4 Turbo. We evaluate each condition with respect to tasks and prompt engineering techniques with 50 repetitions for GPT 3.5 Turbo and 20 repetitions for GPT 4 Turbo due to budgetary considerations. For both model variants and all tasks, we use a temperature setting of $\tau=0$ to elicit factual responses. All other model parameters are set to the default values suggested by OpenAI.
@@ IV-B Prompt Engineering Techniques
Our simulation procedure allows the robot to execute tasks immediately after providing the LLM with the task description along with the function definitions in an initial prompt. Task completion then proceeds as a dialog of function calls generated by the LLM and corresponding function responses generated by the simulation. This approach serves as a baseline for investigating and comparing the effects of various prompt engineering techniques on task completion and time required. We focus on a selection of popular and applicable prompt engineering techniques for task planning and execution in the service robotics domain. Figure 2 gives an overview of this selection.
@@ IV-B2 Chain-of-Thought (CoT)
We integrate CoT [CoT] prompting into our framework by asking the LLM to first provide a textual step-by-step plan detailing how it intends to complete the task given the available functions from the function library. Only after such a plan has been generated, we ask the LLM to call the appropriate functions based on that plan. If the task execution turns out to require many steps, we repeat this planning step every 15 function calls, to ensure that the LLM is always operating according to a plan in case the initial plan is not detailed enough or incomplete.
@@ IV-B4 Example in Prompt
success rate mean time [s] success time [s] success time [s] success time [s]
Baseline 1.00 11.57 0.90 14.72 0.85 23.65 0.45 37.38
AF 1.00 10.69 0.65 12.06 0.60 22.68 0.90 33.64
AF + EiP 1.00 11.21 1.00 16.03 0.70 21.57 0.65 31.97
AF + CoT 1.00 15.14 0.70 18.95 0.70 33.03 0.20 42.32
AF + CoT + EiP 1.00 15.07 1.00 20.69 0.65 26.26 0.80 36.89
AF + ReAct + EiP 1.00 31.05 1.00 31.44 0.95 66.45 0.70 64.25
AF + StD 1.00 10.45 0.90 12.63 0.65 32.05 0.90 29.57
AF + CoT + EiP + StD 1.00 15.52 1.00 20.63 0.80 28.80 0.75 36.44
AF + ReAct + EiP + StD 1.00 29.38 1.00 32.86 1.00 72.34 0.50 72.25
@@ V Results
We find that more complex prompt engineering techniques do not necessarily lead to higher success rates. On the Equals and Distribute tasks with GPT 3.5 Turbo, the techniques without state descriptions outperform the corresponding technique with state descriptions. However, state descriptions do not always have a negative impact. For the GPT 3.5 Turbo Conditional task, the addition of state descriptions (AF + ReAct + EiP + StD) achieves the highest success rate of 72%. We also observe that using only adaptive functions achieves the best success rate with GPT 4 Turbo for the Distribute task, while achieving the worst success rate for the Conditional and Equals tasks. There is also no clear favorite between CoT and ReAct, as the success rates vary depending on the task, model variant, and other prompt engineering techniques used in conjunction.
Due to the additional prompts used in all prompting techniques besides the adaptive functions, the measured task completion time depends on the technique used. In general, we observe that CoT and even more so ReAct show a longer runtime due to the relatively large number of tokens generated by the LLM for planning and reasoning compared to function calls. Using state descriptions increases the time even more for the same reason. Even though the adaptive functions do not utilize additional prompts, we can see a slight reduction in time compared to the baseline, because the LLM is less likely to call functions that result in failures. In a real application, the response time of an LLM can significantly increase the execution time of a task, although functions that require considerable time themselves provide leeway to mask this problem.